OpenStudy (minatsukisaya95):

how do we integrate this xe^(x^2)

OpenStudy (anonymous):

\(u=x^2,\ du=xdx\)

OpenStudy (anonymous):

\[ xe^{x^2}dx\to e^udu \]

OpenStudy (callisto):

A little amendment from above: du = 2x dx

OpenStudy (minatsukisaya95):

isn't du=2x how did it become like that

OpenStudy (minatsukisaya95):

oh ok and then how do we solve it from that

OpenStudy (minatsukisaya95):

sorry i haven't learn how to integrate this yet so i'm clueless bout this question i only have the basic

OpenStudy (anonymous):

You can't guess what the anti-derivative of \(e^u\) is?

OpenStudy (callisto):

\[\int xe^{x^2}dx\]\[=\frac{1}{2}\int 2xe^{x^2}dx\]\[=\frac{1}{2}\int e^{x^2}(2xdx)\] Recall: We let u = x^2, and we know du = 2xdx Can you change the variable in the integral?

OpenStudy (minatsukisaya95):

alright so the answer would be e^(x^2) is that it?

OpenStudy (callisto):

No

OpenStudy (minatsukisaya95):

and then how

OpenStudy (callisto):

What would you get for \(\frac{1}{2}\int e^{x^2}(2xdx)\) when you change the variable from x to u?

OpenStudy (minatsukisaya95):

oh ok i get 1/2(e^(x^2))

OpenStudy (callisto):

You missed something, a constant.

OpenStudy (minatsukisaya95):

is there any other constant other than 1/2

OpenStudy (callisto):

You need to add an arbitrary constant c, or k, whenever you find the indefinite integral.

OpenStudy (callisto):

1/2 is, in fact, the coefficient of e^(x^2)

OpenStudy (minatsukisaya95):

ok i got it so the answer is 1/2(e^(x^2)) + c isn't it correct?

OpenStudy (callisto):

It is correct.

OpenStudy (minatsukisaya95):

yeah!!! thank u u help me a lot.

OpenStudy (callisto):

You're welcome.